# OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation

### v1

@@ 4.1 Experimental Setup
Datasets. CheXpert Plus [8] is a recent large-scale paired image–text resource that links chest X-ray DICOMs to fully de-identified radiology reports, organized into report subsections and annotated for 14 target pathologies (i.e. disease labels). The public release includes $\sim$223K image–report pairs across $\sim$187.7k studies and $\sim$64.7k patients, totaling $\sim$36M report tokens. For data-efficient reinforcement fine-tuning, we subsample a pool of 1,000 studies with a balanced ground-truth label set via CheXpert labeller [26]. We also exclude reports with missing sections. Following recent works [62], we report results on the test split and extract labels from both generated and GT reports by the CheXpert labeller.
Baselines. Following the recent CheXpert Plus benchmark study [62], we compare our proposed approach against a broad suite of methods: 22 task-specific RRG models spanning diverse architectures: ORGan [25], M2KT [70], ASGMD [68], Token-Mixer [71], PromptMRG [27], R2GenRL [51], XProNet [57], MSAT [63], TIMER [66], CvT2DistilGPT2 [46], R2Gen [13], R2GenCMN [14], Zhu et al. [81], CAMANet [58], R2GenGPT [65], WCL [69], VLCI [12], Wang et al. [61], R2GenCSR [60], and the recent MambaXray-B/L [62], as well as CheXagent [15], and 16 open-source LLMs/VLMs, including Vicuna-7B [78], Qwen1.5-7B and Qwen2-7B-Instruct [3], InternLM-7B [7], Llama-2 (7B/13B) [56], Llama-3/3.1 (8B) [18], GPT-2 Medium [52], Orca-2 (7B/13B) [43], DeepSeek-LLM-7B-Chat [5], Yi-1.5 (6B/9B) [73], InternVL-2 [7], and MiniCPM-V2.5 [72]. All task-specific algorithms are fine-tuned on at least the full CheXpert Plus training set, and some of them on multi-source corpora. Thus, our 1K training set represents $\sim$0.45% of the standard 223K split; and only $\sim$0.079% of the 1.27M used by MambaXray [12], which is the current SOTA model.
@@
We fine-tune a 3B Qwen-2.5-VL policy [4] on the 1,000 CheXpert Plus studies in an RL-only setting (no SFT/alignment), using Hugging Face TRL’s GRPOTrainer with transformers and accelerate. Training runs on four NVIDIA A10 GPUs (24 GB) with PyTorch 2.7.1/CUDA 12.1. We empirically set batch size $B=16$. For each prompt, we sample $K=8$ completions, compute the FactS reward per completion, and optimize the objective in Eq. (11): GRPO provides the default learning signal, while DPO is invoked adaptively via the ZRR schedule. To prioritize GRPO yet correct hard, zero-reward cases, we map the EMA zero-reward rate $\tilde{z}_{i}^{(t)}$ with $\alpha=0.5$ in Eq. (9) to a bounded mixing weight $w_{i}^{(t)}$ in Eq. (10) with $w_{\min}\!=\!0.05,\quad w_{\max}\!=\!0.15,\quad\gamma\!=\!2.$ Thus, $w_{i}^{(t)}\!\to\!w_{\max}$ only when a prompt’s ZRR approaches 1 (no usable GRPO signal), and it quickly decays toward $w_{\min}$ once GRPO yields non-zero rewards, ensuring GRPO dominates ordinary updates and DPO remains a light, targeted aid. We adopt two length-aware variants to handle variable-length reports: DR.GRPO [38] to mitigate length bias in on-policy updates, and LN-DPO [2] to normalize preference margins by sequence length.
@@ 4.2 Main Results and Analysis on CheXpert Plus
As shown in Tab. 1, OraPO+FactS sets the new SOTA F1 = 0.341 (best) and Recall = 0.832 (best) on CheXpert Plus, significantly surpassing best-performing baselines by $\mathbf{+160.8\%}$ in recall. When compared with the most recent baseline, VLCI, our proposed method outperforms VLCI by $\mathbf{+109.2\%}$ in F1 and $\mathbf{+375.4\%}$ in recall. While our precision trails MambaXray-L, this is a deliberate trade-off between emphasizing per-finding coverage and accurate statements of detected findings. In practice, recall-oriented drafts are preferred, as radiologists perform the final review, they often prefer high-sensitivity assistance [23]. Overall, our method sets the new SOTA performance in F1 and recall, with three orders of magnitude less training data.
Against Alignment-based SFT for LLM/VLM. In this experiment, we compare our framework against R2GenGPT [65], a popular alignment-based SFT approach that is widely used for training large LLM/VLM. In this approach, a visual encoder is aligned with a language model into the same space via supervised fine-tuning on paired image-report data. We report the performance of R2GenGPT [65] with different LLM/VLM backbones in Tab. 2. Our approach is trained with a compact 3B policy on 1K CheXpert Plus samples. Under this small-data setup, OraPO attains $\mathbf{F1=0.341}$ and $\mathbf{Recall=0.832}$, significantly outperforming larger aligned backbones trained on 223K: F1 improves over InternLM by $\mathbf{+20.1\%}$ and over Qwen-2 (0.261) by $\mathbf{+30.7\%}$; recall improves over InternLM by $\mathbf{+203.6\%}$ and over GPT-2 Medium by $\mathbf{+347.3\%}$. These results validate that our proposed OraPO with fact-level supervision is a superior learning paradigm than standard vision–language alignment, achieving higher performance with less than half backbone parameters and markedly less training data.
@@ 4.3 Ablation Study
To study the effectiveness and efficiency of our proposed method, we evaluate four model variants on CheXpert Plus using macro Precision, Recall, and F1. As reported in Tab. 3, the second row is a naïve GRPO-based RLHF baseline that uses accuracy as reward (as in DeepSeek-R1 [21]), which yields only marginal gains over the base model without any domain-specific training (row 1). Adding FactS to GRPO (row 3) lifts F1 from 0.089 to 0.291 $(+227\%)$, Recall from 0.162 to 0.605 (+274%), and Precision from 0.026 to 0.204 (+685%), showing that fact-level entailment provides dense, clinically aligned supervision. Layering OraPO on top (row 5) further improves F1 to 0.341 (+17.2% vs. FactS-only), with Recall improved by (+37.5%) and Precision improved by (+16.2%). Notably, OraPO+FactS with only 400 studies (row 4) attains F1 == 0.296, Precision == 0.217 and Recall == 0.732—–surpassing the FactS-only 1K setting (F1 == 0.291, Precision == 0.204, and Recall == 0.605)–—providing evidence of OraPO’s superb data efficiency. Overall, FactS supplies faithful, sentence-level rewards, while OraPO converts zero-reward groups into preference updates, yielding higher precision & recall and earlier learning on scarce yet clinically important signals.
@@ 5 Conclusion
In this work, we present OraPO, an oracle-educated RL training algorithm for high-quality radiology report generation under tight data and compute budgets. To our knowledge, we are the first to couple DPO with GRPO: when GRPO exploration fails, we recycle those rollouts as negatives and apply a lightweight DPO step to recover useful gradients. In parallel, we design the FactS reward to provide dense, fact-level, clinically faithful supervision tailored to RRG. Comprehensive experiments show OraPO setting the new SOTA performance while using 2–3 orders of magnitude fewer training samples.
We have identified a number of promising future work directions. (i) While designed for the RRG task, OraPO can be applied to other hard, general-domain tasks where vanilla GRPO struggles with sparse success signals (e.g., multi-step mathematical reasoning and code synthesis). We will extend experiments to other tasks to validate the effectiveness of OraPO. (ii) Due to compute constraints, our results currently are limited to a 3B backbone VLM with only 1K studies for training. We plan to study OraPO’s scaling behavior across model sizes and data/training budgets. (iii) While our reward design already bridges the gap between verifiable and non-verifiable model outputs by leveraging LLMs as a judge, it still falls short of handling genuinely non-verifiable rewards. Inspired by the recent Rubrics as Rewards technique [20], we plan to extend our methodology to address the challenge of non-verifiable reward modeling. (iv) FactS mainly scores disease presence/absence and underweights attributes (location, laterality, size/severity, temporal change). We will design a more fine-grained fact-to-attribute entailment reward for addressing the misalignment issue on diagnostic attributes.

### v2

@@ 4.1 Experimental Setup
Datasets. CheXpert Plus [8] is a recent large-scale paired image–text resource that links chest X-ray DICOMs to fully de-identified radiology reports, organized into report subsections and annotated for 14 target pathologies (i.e. disease labels). The public release includes $\sim$223K image–report pairs across $\sim$187.7K studies and $\sim$64.7k patients, totaling $\sim$36M report tokens. MIMIC-CXR [41] is a large, publicly available paired image–text resource of chest X-rays in DICOM format with contemporaneous, de-identified radiology reports. It contains 377,110 images from 227,943 imaging studies covering 65,379 patients. Each study includes one or more views (typically frontal and/or lateral) and a free-text report that often follows Findings and Impression sections. For data-efficient fine-tuning, we sample 1,000 studies per dataset (i.e. 1,000 from CheXpert-Plus and 1,000 from MIMIC-CXR) with a balanced ground-truth label set via CheXpert labeller [36]. We also exclude reports with missing sections. Following recent works [92], we report results on the test split and extract labels from both generated and GT reports by the CheXpert labeller. We further evaluate on the CheXpert Plus human gold set, whose expert-verified labels provide a more reliable basis for evaluation and stronger conclusions.
Baselines. Following the recent benchmarking study [92], we compare our proposed approach against a broad suite of methods: 28 task-specific RRG models spanning diverse architectures and 16 open-source LLMs/VLMs, as listed in Tables 1, 2 and 3. All task-specific algorithms are fine-tuned on at least the full training set of the dataset used in each experiment (CheXpert Plus or MIMIC-CXR), with some methods additionally incorporating multi-source corpora. Thus, our 1K training set sampled from CheXpert Plus, represents $\sim$0.45% of the standard 223K split; and only $\sim$0.079% of the 1.27M used by MambaXray [12], which is the current SOTA model. A similar ratio applies to MIMIC-CXR.
@@
We fine-tune a 3B Qwen-2.5-VL policy [4] on the 1,000 CheXpert Plus studies in an RL-only setting (no SFT/alignment), using Hugging Face TRL’s GRPOTrainer with transformers and accelerate. Training runs on four NVIDIA A10 GPUs (24 GB) with PyTorch 2.7.1/CUDA 12.1. We empirically set batch size $B=16$. For each prompt, we sample $K=8$ completions, compute the FactS reward per completion, and optimize the objective in Eq. (11). More detailed description of the hyper-parameters for OraPO can be found in the Supplementary.
@@ 4.2 Main Results and Analysis on CheXpert Plus
As shown in Table 1, OraPO+FactS sets the new SOTA F1 = 0.341 (best) and Recall = 0.832 (best) on CheXpert Plus, significantly surpassing best-performing baselines by $\mathbf{+160.8\%}$ in recall. When compared with the most recent baseline, VLCI, our proposed method outperforms VLCI by $\mathbf{+109.2\%}$ in F1 and $\mathbf{+375.4\%}$ in recall. While our precision trails MambaXray-L, this is a deliberate trade-off between emphasizing per-finding coverage and accurate statements of detected findings. In practice, recall-oriented drafts are preferred. Since radiologists perform the final review, they tend to favour AI assistance that has higher sensitivity (even at the cost of precision) [31]. Overall, our method sets the new SOTA performance in F1 and recall, with three orders of magnitude less training data.
Against Alignment-based SFT for LLM/VLM. In this experiment, we compare our framework against R2GenGPT [94], a popular alignment-based SFT approach that is widely used for training large LLM/VLM. We report the performance of R2GenGPT [94] with different LLM/VLM backbones in Table 2. Our approach is trained with a compact 3B policy on 1K CheXpert Plus samples. Under this small-data setup, OraPO attains $\mathbf{F1=0.341}$ and $\mathbf{Recall=0.832}$, significantly outperforming larger aligned backbones trained on 223K. These results validate that our proposed OraPO with fact-level supervision is a superior learning paradigm to standard vision–language alignment, achieving higher performance with less than half backbone parameters and markedly less training data.
Evaluation on Human Gold Label Set. To obtain a more faithful evaluation, we assess models on the CheXpert validation set with human gold labels: 200 studies from 200 patients independently annotated by three board-certified radiologists, where a majority vote defines the ground truth [36]. Table 5 reports macro Precision/Recall/F1 across the 14 findings. With only 1K training cases, our method reaches 0.234/0.641/0.288, surpassing MambaXray-B, MambaXray-L and CheXagent in both Recall and F1. These gains arrive while using 0.1% of MambaXray’s data and around 0.01% of CheXagent’s.
We also evaluate two commercial APIs (GPT-4.1 [70] and GPT-5 Thinking [71]) on the CheXpert validation set with human gold labels, using the identical image input and report-generation prompt. GPT-5 is configured with a medium reasoning level. As reported in Table 5, our model outperforms GPT-4.1 by +110.2% in recall and +13.8% in F1. Relative to GPT-5 Thinking, our model achieves +52.3% higher recall, though F1 is -4.3% lower. Critically, compared to GPT-5 Thinking, our compact 3B model delivers real-time inference (3.3 s/image vs. 25.2 s/image), a lightweight and transparent footprint (3B vs. trillions of parameters), and cheaper deployment that eliminates per-token API fees. These advantages make our method substantially more cost-efficient and suitable for time-critical clinical use than calling commercial APIs. Overall, the results confirm that our improvements persist when measured against expert gold labels, while being dramatically more data-efficient.
4.3 Main Results and Analysis on MIMIC-CXRs
On MIMIC-CXR, our approach attains F1 = 0.357 and Recall = 0.891 using just 1K training samples, as shown in Table 3. This yields a +5.0% F1 gain over the best baseline (0.340, MambaXray-L) and a +153.8% recall gain over the strongest baseline recall (0.351, CXRMate). The markedly higher recall under macro averaging indicates substantially fewer omissions across both common and rare pathologies, aligning with clinical preference for high sensitivity and omission avoidance [62, 78, 76, 116, 69]. At the same time, the overall F1 is the best among all baselines, despite a precision of 0.242. Together with the CheXpert-Plus results, these findings show that OraPO generalises across datasets and achieves state-of-the-art clinical effectiveness with 2–3 orders-of-magnitude less data than prior approaches.
@@ 4.4 Ablation Study
To study the effectiveness and efficiency of our proposed method, we evaluate four model variants on CheXpert Plus using macro Precision, Recall, and F1. As reported in Table 4, the second row is a naïve GRPO-based RLHF baseline that uses accuracy as reward (as in DeepSeek-R1 [29]), which yields only marginal gains over the base model without any domain-specific training (row 1). Adding FactS to GRPO (row 3) lifts F1 from 0.089 to 0.291 $(+227\%)$, Recall from 0.162 to 0.605 (+274%), and Precision from 0.026 to 0.204 (+685%), showing that fact-level entailment provides dense, clinically aligned supervision. Layering OraPO on top (row 6) further improves F1 to 0.341 (+17.2% vs. FactS-only), with Recall improved by (+37.5%) and Precision improved by (+16.2%). Notably, OraPO+FactS with only 400 studies (row 4) attains F1 == 0.296, Precision == 0.217 and Recall == 0.732—–surpassing the FactS-only 1K setting (F1 == 0.291, Precision == 0.204, and Recall == 0.605)–—providing evidence of OraPO’s superb data efficiency. We further observe that the GRPO+SFT variant (with FactS) collapses to very low recall (0.176) and F1 (0.106), a -70.9% and -63.6% drop vs. FactS+GRPO, respectively. The core issue here is that plain SFT only imitates ground-truth prose: the model learns “what to say when it’s right,” but not “what not to say,” so reports that omit all true positives aren’t explicitly discouraged, leading to recall and F1 collapses. Beyond that, SFT on small datasets is known to be unstable and sample-inefficient: fine-tuning with limited data often over-fits, degrading generalisation [73, 22]. OraPO fixes this by explicitly pushing probability mass away from “no-TP” outputs such that the recall is effectively boosted when trained with limited set. Overall, FactS supplies faithful, sentence-level rewards, while OraPO converts zero-reward groups into preference updates, yielding higher precision and recall and faster convergence on scarce yet clinically important signals.
@@ 5 Conclusion
In this work, we present OraPO, an oracle-educated RL training algorithm for high-quality radiology report generation under tight data and compute budgets. To our knowledge, we are the first to couple DPO with GRPO: when GRPO exploration fails, we recycle those rollouts as negatives and apply a lightweight DPO step to recover useful gradients. In parallel, we design the FactS reward to provide dense, fact-level, clinically faithful supervision tailored to RRG. Comprehensive experiments show that OraPO sets the new SOTA performance while using 2–3 orders of magnitude fewer training samples. In future work, we target four directions: (1) Extend OraPO beyond RRG to sparse-signal tasks (e.g. multi-step math, code) where vanilla GRPO struggles. (2) Study scaling across model size and data/training budgets beyond our current 3B/1K setup. (3) Tackle non-verifiable rewards via rubric-driven signals (e.g. Rubrics as Rewards [28]). (4) Enrich FactS with fine-grained attribute entailment (location, laterality, size/severity, temporal change) to reduce diagnostic attribute misalignment.
Supplementary Material of OraPO: Oracle-educated Reinforcement Learning
for Data-efficient and Factual Radiology Report Generation
This supplementary material includes:
  • Algorithm 1: A detailed, step-by-step description of the proposed method.
  • Section A: Comprehensive implementation details.
  • Section B: Additional experimental results and analysis on the MIMIC-CXR dataset.
  • Section C: Qualitative studies, including examples of generated reports and assessments on extracted facts.
Appendix A More Implementation Details
[…]
Table 6: Hyperparameter settings of the proposed method (selected in underline).
Parameter Description Search Range / Setting
$B$ Effective Batch Size $\{\underline{16}\}$
LR Learning Rate $\{1e-6,\underline{2.5e-7}\}$
$K$ GRPO sampled group size $\{4,\underline{8},16\}$
$\alpha$ EMA momentum for zero-reward rate (ZRR) $\{0.4,\underline{0.5},0.6\}$
$w_{\min}$ Minimum DPO mixing weight $\{0.02,\underline{0.05},0.1\}$
$w_{\max}$ Maximum DPO mixing weight $\{\underline{0.15},0.3\}$
$\gamma$ Sharpening exponent for mapping ZRR to $w_{i}^{(t)}$ $\{1.0,\underline{2.0}\}$
Table 7: Experimental results (micro averaging) on the MIMIC-CXR dataset [8].
Algorithm Venue Precision, Recall, F1 Train Size
MedRAT [32] ECCV24 0.285, 0.265, 0.227 223K
MET [93] CVPR23 0.364, 0.309, 0.311 223K
KiUT [35] CVPR23 0.371, 0.318, 0.321 369K
DCL [46] CVPR23 0.471, 0.352, 0.373 223K
RGRG [83] CVPR23 0.524, 0.474, 0.498 223K
CoFE [47] ECCV24 0.489, 0.370, 0.405 223K
COMG [27] WACV24 0.424, 0.291, 0.345 223K
MPO [99] AAAI25 0.436, 0.376, 0.353 223K
MAN [80] AAAI24 0.411, 0.398, 0.389 223K
Med-LLM [55] MM24 0.412, 0.373, 0.395 223K
B-LLM [52] AAAI24 0.465, 0.482, 0.473 223K
EKAGen [6] CVPR24 0.517, 0.483, 0.499 223K
MambaXray-L [92] CVPR25 0.561, 0.460, 0.505 1.27M
MLRG [53] CVPR25 0.549, 0.468, 0.505 240K
Ours 0.342, 0.811, 0.481 1K
Algorithm 1 The proposed OraPO with FastS Reward for Radiology Report Generation
Input initial policy model $\pi_{\theta}^{\text{init}}$; dataset $\mathcal{D}$ with studies $(x,y^{\star})$ and prompts $p$; label set $L$; group size $K$; hyperparameters $\varepsilon,\lambda_{\mathrm{KL}},\tau,\alpha,\gamma,w_{\min},w_{\max},\beta,\xi$; outer loops $I$, inner steps $M$.
Output $\pi_{\theta}$
1:policy model $\pi_{\theta}\leftarrow\pi_{\theta}^{\text{init}}$
2:for $\text{iteration}=1,\ldots,I$ do
3:  reference model $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$
4:  for $\text{step}=1,\ldots,M$ do
5:   Sample a minibatch $\{(x_{i},p_{i},y_{i}^{\star},z_{i}^{\star})\}^{B}\subset\mathcal{D}$
6:   Update the behaviour policy $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
7:   Sample $K$ reports $\{\hat{y}_{i,j}\}_{j=1}^{K}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x_{i},p_{i})$ for each input radiology image $x_{i}$ in the batch
8:   Compute FactS reward: $\{r(x_{i},\hat{y}_{i,j})\}_{j=1}^{K}$ for each sampled report $\hat{y}_{i,j}$;
9:   for $\text{OraPO iterations}=1,\ldots,\mu$ do
10:     Compute GRPO objective: $L_{\mathrm{GRPO}}(x_{i},p_{i})$;
11:     Compute DPO objective: $L_{\mathrm{DPO}}(x_{i},p_{i})$;
12:     Compute OraPO objective: $\mathcal{L}_{\mathrm{OraPO}}$ by mixing $L_{\mathrm{DPO}}(x_{i},p_{i})$ and $L_{\mathrm{GRPO}}(x_{i},p_{i})$;
13:     Update the policy model $\pi_{\theta}$ by minimizing $L_{\mathrm{OraPO}}(x_{i},p_{i})$.
14:   end for
15:  end for
16:end for
Appendix B More experimental results and analysis on MIMIC-CXR
Table 7 complements our main macro-averaged results by reporting micro-averaged Precision, Recall, and F1 on MIMIC-CXR. Macro averaging, used in the main paper, treats each disease label equally and highlights performance on rare findings. Micro averaging instead weights labels by their frequency, answering a different question: across all positive label instances in the dataset, how many do we miss. This view is especially focusing on common conditions which dominate the label distribution.
Under this micro-averaging metric, OraPO remains highly competitive while using 2–3 orders of magnitude less data. With only 1 K training reports, OraPO attains a micro recall of 0.811 and micro F1 of 0.481. The best baseline recall is EKAGen at 0.483, so OraPO improves micro recall by 67.9% relative to this strong model. Clinically, such high recall means that, across all truly abnormal cases, our method misses far fewer findings, which is critical because false negatives on common diseases are usually more harmful than extra false positives.
Although our method achieves the highest recall, its micro F1 remains close to that of the strongest fully supervised approaches. The best baseline F1 is 0.505 (MambaXray-L and MLRG), whereas OraPO reaches 0.481, only a gap of 0.024. This trade-off is achieved with extreme data efficiency: compared with methods trained on the full 223K MIMIC-CXR split, OraPO uses 99.6% fewer labeled reports, and compared with MambaXray-L’s 1.27M training pairs, about 99.9% fewer. Overall, the micro results show that OraPO delivers clinically preferred high recall with competitive F1, while training with only 1K samples.
Appendix C Qualitative Study
In Figure 3, we present three examples that demonstrate how OraPO with FactS reward generates clinically accurate reports while maintaining factual grounding through atomic fact extraction and entailment checking.
Example 1 showcases ideal performance on a challenging case with cardiomegaly and edema. The model produces a clinically coherent narrative that captures the typical imaging pattern of cardiac-related pulmonary edema. Generated facts such as “cardiac silhouette is enlarged” and “diffuse interstitial opacities seen centrally” directly map to the ground-truth labels, with no false positives.
Example 2 highlights OraPO’s handling of complex cases with five ground-truth labels. The generated report correctly identifies all target pathologies, including the challenging bilateral nodular pulmonary lesions. While the model predicts pleural effusion absent from the ground-truth labels, the generated facts (“pleural effusions present” with “decreased interval changes”) suggest a false positive predicted by the model.
Example 3 illustrates robust multi-pathology detection spanning four conditions: lung opacity, consolidation, pneumonia, and pleural effusion. The model generates anatomically precise descriptions (“left perihilar opacity with patchy distribution”) that provide localization detail often absent in baseline systems. All ground-truth labels receive explicit supporting facts, demonstrating the FactS reward’s effectiveness in maintaining high recall across co-occurring pathologies. The model correctly attributes findings to specific lung regions, which aids radiologist review.
Across these examples, OraPO consistently produces factually grounded reports where each diagnostic statement traces back to verifiable atomic facts that entail specific labels. This stands in contrast to fluency-optimized baselines that generate plausible prose without guaranteed factual alignment. The FactS reward’s sentence-level supervision enables the model to maintain clinical accuracy even when handling rare pathologies or multi-label complexity, directly addressing the recall-precision trade-off critical in clinical deployment.
Here’s a concise qualitative analysis paragraph you can place right below Fig. 3:
C.1 Qualitative NLG Metric Analysis
Although the ROUGE-L F1 values in Figure 3 are modest, they largely reflect surface mismatch rather than clinical errors: OraPO is not trained to imitate the phrasing or template of GT reports. Instead of imitating the GT writing style, our method directly optimizes a fact-level reward. The model first extracts atomic clinical statements, and then checks whether these statements are entailed by the ground-truth labels. This training objective naturally prioritizes clinical correctness and verifiability, not surface-level similarity. In contrast, reference-overlap metrics such as ROUGE mainly measure wording overlap. They reward matched phrasing but do not penalize factual mistakes, missing findings, or even cross-sentence contradictions in long reports. Concretely, our FactScore-based reward treats the report as its own rationale, maps sentences to label-level predictions, and yields dense, interpretable feedback to the policy. In the three qualitative cases, the generated statements are clinically grounded and aligned with the GT labels: Example 1 captures cardiomegaly and edema with supporting facts; Example 2 correctly identifies all target pathologies (with one flagged false-positive effusion); and Example 3 localizes a left perihilar opacity and supports consolidation/pneumonia/effusion—all consistent with the label set despite low lexical overlap with GT wording. This behavior is consistent with our quantitative results. OraPO is deliberately recall-oriented, which is clinically the safer operating point. On CheXpert Plus, it achieves state-of-the-art label-level recall and F1 while using only about 0.1% of the usual training data, and it also attains very high micro-recall. Therefore, low n-gram overlap with the reference should not be interpreted as low clinical faithfulness.
Image GT Reports GT Labels Our Generated Reports Predicted Labels and Supporting Facts ROUGE-L F1
Refer to caption Single frontal view of the chest demonstrates indistinct pulmonary vasculature bilaterally, likely representing mild pulmonary edema. Mild cardiomegaly is seen, unchanged from prior exam. No soft tissue or bony abnormalities are seen. Cardiomegaly, Edema Ribs: The thoracic rib contour is smooth and complete. There is mild opacity seen in the mid to left hemithorax. Pectoral Girdle: Both clavicles are seen in their normal positions. Diaphragm: The left diaphragm appears increased in curvature without significant deviation. Aortic Arch/Trunk: The aortic arch appears well defined. Cardiomediastinum: There are diffuse interstitial opacities seen centrally on the image, consistent with pulmonary edema. The right border of the heart is poorly defined and the cardiac silhouette is enlarged. The pulmonary vascularity is increased centrally but diffuse through the lung fields and the cardiac silhouette appears to be expanding to the right of the sternum. Pulmonary Vasculature: The pulmonary vasculature appears increased centrally. Lung Opacity: There is diffuse sub-pleural opacity centrally, with extension towards the left hemithorax. Increased pulmonary vasculature seen centrally, consistent with pulmonary edema. No other areas of consolidation seen. This appearance is consistent with pulmonary edema. There are small sub-pleural opacities seen centrally. - Cardiomegaly: ’Cardiac silhouette is enlarged’ and ’Cardiac silhouette appears to be expanding to the right of the sternum’ both indicate cardiomegaly. - Edema: ’Diffuse interstitial opacities seen centrally’, ’Pulmonary vascularity is increased centrally’, ’Pulmonary vascularity is diffuse through the lung fields’, and ’Appearance is consistent with pulmonary edema’ all indicate edema.” 0.102
Refer to caption Within the right middle lobe there is a large consolidation, consistent with pneumonia, which obscures the right heart border. A small left upper lobe calcified granuloma is present. There is mild calcification of the aortic arch. No pleural effusions. Mild cardiomegaly. No bony abnormalities are appreciated. Cardiomegaly, Lung Opacity, Lung Lesion, Consolidation, Pneumonia Cardiomediastinal silhouette: mildly increased in size; cardiac contour is flattened Nodular pulmonary lesions present bilaterally - no growth or interval increase Pleural effusions present - decreased interval changes Pneumonic opacities present, right mid and left lobe, bilateral, but limited in extent Right diaphragmatic elevation - Cardiomegaly: ’Cardiomediastinal silhouette is mildly increased in size’ directly supports presence of cardiomegaly. - Lung Opacity: Facts ’Pneumonic opacities present in right mid lobe’, ’Pneumonic opacities present in left lobe’, and ’Nodular pulmonary lesions present bilaterally’ all describe opacities. - Lung Lesion: ’Nodular pulmonary lesions present bilaterally’ directly supports presence of lung lesions. - Consolidation: Facts ’Pneumonic opacities present in right mid lobe’, ’Pneumonic opacities present in left lobe’, and ’Pneumonic opacities are bilateral’ support presence of consolidation. - Pneumonia: Facts ’Pneumonic opacities present in right mid lobe’, ’Pneumonic opacities present in left lobe’, and ’Pneumonic opacities are bilateral’ support presence of pneumonia. - Pleural Effusion: ’Pleural effusions present’ directly supports presence of pleural effusion. 0.114
Refer to caption Initial chest radiograph at 1130 hours demonstrates left lower lobe focal air space opacity which could represent pneumonia. The remainder of the lungs are clear. The cardiomediastinal silhouette is normal. There is blunting of the left costophrenic angle which could represent a small pleural effusion. Follow-up radiograph 1530 hours demonstrates no interval change. Lung Opacity, Consolidation, Pneumonia, Pleural Effusion Single ventricular image. Cardiac silhouette is no grossly enlarged (LVED: 55-57 mm, LVED: 55-57 mm) and changes relative to previous films. No evidence of marked cardiomegaly in the presence of mild respiratory straining. Left perihilar opacity with patchy distribution. Small pleural effusion present, mild volume and interval increase noted.Cardiac silhouette interval increase noted. - Lung Opacity: Fact ’Left perihilar opacity with patchy distribution’ describes an opacity, confirming existence. - Consolidation: Fact ’Left perihilar opacity with patchy distribution’ describes an opacity consistent with consolidation. - Pneumonia: Fact ’Left perihilar opacity with patchy distribution’ describes an opacity that could represent pneumonia. - Pleural Effusion: Facts ’Small pleural effusion present’, ’Mild volume increase in pleural effusion’, and ’Interval increase in pleural effusion’ all confirm pleural effusion. 0.164
Figure 3: X-ray image and its corresponding ground-truth, along with the output of our model generation report on the ChexPert Plus dataset. The mismatch sentence in the reports are highlighted using different colors.